OpenStudy (anonymous):

In the function y-1=(4x)^2+7 , what effect does the number 4 have on the graph, as compared to the graph of y=x^2 ? A. It stretches the graph vertically by a factor of 4. B. It shrinks the graph horizontally to 1/4 the original width. C. It shrinks the graph vertically to 1/4 the original height. D. It stretches the graph horizontally by a factor of 4. ***Idk what this one means :/ please explain :) thanks! :D

jimthompson5910 (jim_thompson5910):

are you sure it's not y-1=4(x)^2+7 or y-1=4x^2+7

OpenStudy (anonymous):

here ill draw it out :)

OpenStudy (anonymous):



jimthompson5910 (jim_thompson5910):

hmm interesting

OpenStudy (anonymous):

yeah idk what it means :/

OpenStudy (anonymous):

do u understand it? :D

jimthompson5910 (jim_thompson5910):

i think it's A, but that would mean it would have to be either y-1=4(x)^2+7 or y-1=4x^2+7 the coefficient for the x^2 term dictates how you stretch/compress the parent function

OpenStudy (anonymous):

is there a difference when u change it like that??

jimthompson5910 (jim_thompson5910):

but the way you drew it, it's way off (you would square 4 to get 16, but that's not an answer choice)

jimthompson5910 (jim_thompson5910):

yes there is

OpenStudy (anonymous):

here is exactly how it looks :)

OpenStudy (anonymous):

idk if that helps? haha

OpenStudy (anonymous):

wait but didn't u say earlier the k part is the up/down?? and k=8 in this equation right?? so would it still be stretching vertically ??? is the effect of the 4 a VERTICAL stretch?

jimthompson5910 (jim_thompson5910):

no, that's shifting vertically

jimthompson5910 (jim_thompson5910):

4 is part of the vertical stretch, but it's really 4^2 = 16

OpenStudy (anonymous):

wait i'm confused.. so how do i find the answer to this??

jimthompson5910 (jim_thompson5910):

there has to be a typo

jimthompson5910 (jim_thompson5910):

my guess is that they meant to say y-1=4x^2+7

OpenStudy (anonymous):

here is something i found :) idk if that helps tho cuz i think u know this already haha

jimthompson5910 (jim_thompson5910):

i do and those aren't used in this particular case

OpenStudy (anonymous):

okay haha is it this that is used?

jimthompson5910 (jim_thompson5910):

oh, ok, i see what they're using

jimthompson5910 (jim_thompson5910):

x^2 becomes (4x)^2 so they replaced 'x' with '4x'

OpenStudy (anonymous):

so yeah? the thingy i just posted is what we are using?

jimthompson5910 (jim_thompson5910):

agreed?